OpenStudy (anonymous):

Essay; show all work. Last fall, a gardener planted 62 iris bulbs. She found that only 48 of the bulbs bloomed in the spring. Find the empirical probability that an iris bulb of this type will bloom. Give answer as a fraction in lowest terms. How many of the bulbs should she plant next fall if she would like at least 53 to bloom?

OpenStudy (mathmate):

Out of 62 trials, 48 were successes, so the empirical probability would be: p = _____ divided by _______.

OpenStudy (anonymous):

62 divided by 48 = 1.291

OpenStudy (mathmate):

One of the axioms of probability says that ALL probabilities must fall between zero and one, inclusivement. That is to say: 0<=p<=1. So whenever you get a proabability great than one, or less than zero, you need to recheck you work. This helps you spot errors. With this in mind, would you like to try again?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

48/62=24/31 . x/53=24/31 proportion x=1272/31 =41.03 more than 41 should be planted.

OpenStudy (mathmate):

24/31 is correct, well done. Now, your equation: x/53=24/31 is a good attempt, because we need a proportion equation. 24/31 is #success/#trials. For x and 53, 53 is the expected number of successes, and x is the number of trials needed. When we put an equal sign, both sides must represent the same thing, in our case, it's #success/#trials, right? Would you like to try the proportion equation again?

OpenStudy (anonymous):

24/31 b. let x = # planted so (24/31)x = 53 she wants to bloom and solve for x gives 69 bulbs. Ans: she should plant 69 bulbs.

OpenStudy (mathmate):

Perfect! Well done!

OpenStudy (anonymous):

thank you :)